OpenStudy (anonymous):

write an equation of a line that passes through (-4,3) and is perpendicular to the line y=7/5x-3/5

OpenStudy (anonymous):

We are looking for a line perpendicular to the one given. Therefore, we need to find the negative inverse of the slope of the given line. 7/5 --> -5/7 Now, we simply substitute in our new slope and the point given to find the y-int. Then we will be able to write an equation of the perpendicular line. 3 = -5/7(-4) + b 3 = 20/7 + b 21/7 = 20/7 + b 1/7 = b Thus, the line is: y = -5/7x + 1/7

OpenStudy (anonymous):

these were the options A.5x+7y=1 B. 7x+5y=-13 C.7x+5y=-1 D.9x+5y=7

OpenStudy (anonymous):

Ahh, they wanted the standard form of the line. Simply move your variable terms over to one side: y = -5/7x + 1/7 5/7x + y = 1/7 Now get rid of fractions by multiplying all terms by 7: A) 5x + 7y = 1